A ladder 13 ft long leans against a wall. The bottom of the ladder is 5 ft from the wall. How much would the lower end of the ladder have to be pulled away so that the top end would be pulled down by 2 ft?
Answer by kapilsinghi(68)

Just see the ladder against wall as a right angled triangle
Length of Ladder = 13m
Bottem of ladder = 5m
So Top end of ladder = Squr(13*13 - 5*5)
= Squr(169 - 25)
= Squr(144)
= 12
If top end is to be pulled down by 2 m then
Length of ladder = 13 m
top Of ladder = 12 - 2 =10m
bottem of ladder = Squr(13*13 - 10*10)
= Squr(169 - 100)
= Squr(69)
= 8.3m
So bottom Of ladder is to be shifted by 8.3m - 5m = 3.3m
kapilsinghi123@gmail.com